1. Maria can read 20 pages of economics in an hour. She can also read 50 pages of sociology in an hour. She spends 5 hours per day studying.
  2. Draw Maria’s production possibilities frontier
  3. for reading economics and sociology.
  4. What is Maria’s opportunity cost of reading
  5. 100 pages of sociology? 
1
Expert's answer
2022-02-14T15:30:55-0500

"\text{1.}\, \text{E}=20\ \text{pages}\\

\text{S}=50\ \text{pages}\\

\text{If Maria spends 5hours on reading Economics}\\

\text{she would read} (20*5)\text{pages}=100\text{pages}\\

\text{If she spends 5hours on Sociology,}\\

\text{she would read} (50*5)\text{pages}=250\text{pages}\\

\text{if she decides to spend 2.5hours each on them,}\\

\text{she would read};\\

(20*2.5)\text{pages}=50\text{pages of Economics}\\

(50*2.5)\text{pages}=125\text{pages of Sociology}\\"

"\text{2.}\, \text{To read 1 page of Sociology, she is giving up,}\\
\dfrac{50pages}{20pages}=2.5\ \text{pages of Economics}\\
\text{Therefore, to read 100 pages of Economics,}\\
\text{Maria would have to give up}\\
(100*2.5)=250\,\text{pages of Economics}\\
\text{which is her opportunity cost.}"